OpenStudy (anonymous):

Find the difference quotient for the equation f(x)=-x^2+x ?

OpenStudy (anonymous):

Do you have to show work, or can you take a shortcut?

OpenStudy (anonymous):

show work if possible

OpenStudy (anonymous):

Ok, you know what 'difference quotient' means right? It's just like the slope formula for straight lines, but uses differentials. \[DQ=\frac{f(x+\Delta x)-f(x)}{\Delta x}\]

OpenStudy (anonymous):

Do you know how to set it up from there?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

(-(x+h)^2+(x+h))-(-x^2+x))/h = (-(x^2+2xh+h^2)+x+h+x^2-x)/h = (-x^2-2xh-h^2+x+h=x^2-x)/h = (-2xh-h^2+h)/h = -2x+1

OpenStudy (zzr0ck3r):

which is the derviative of -x^2 + x

OpenStudy (zzr0ck3r):

(f(x+h) - f(x))/h

OpenStudy (zzr0ck3r):

hope that helps

OpenStudy (zzr0ck3r):

the first equal sign on the second line should be +

OpenStudy (zzr0ck3r):

(-(x+h)^2+(x+h))-(-x^2+x))/h = (-(x^2+2xh+h^2)+x+h+x^2-x)/h = (-x^2-2xh-h^2+x+h+x^2-x)/h = (-2xh-h^2+h)/h = -2x+1

OpenStudy (zzr0ck3r):

have fun in calculus, its really fun:)

OpenStudy (zzr0ck3r):

does this make sense @jalotaibi ?

OpenStudy (anonymous):

If \[f(x)=-x^2+x\] then the difference quotient, \[\frac{f(x+\Delta x)-f(x)}{\Delta x} = \frac{-(x+\Delta x)^2+(x+\Delta x)-(-x^2+x)}{\Delta x}.\]

OpenStudy (anonymous):

Then just simplify from there.